Find the inverse of the function f(x)= -2log3(x-1)+4
The '3' is the base of the logarithm. The coefficient and the '+4' part are throwing me off so if you could show me how to do it with or without the answer I'd appreciate it

Answer by stanbon(57395)

Find the inverse of the function
f(x)= -2log3(x-1)+4
----
1st: Interchange x and y
x = -2log3(y-1) + 4
----
2nd: Solve for "y":
-2log3(y-1) = x-4
---
divide by -2 to get:
log3(y-1) = (4-x)/2
---
y-1 = 3^(4-x)/2
----
y = 3^[(4-x)/2] +1
----
That is the inverse function.
==============================
Cheers,
Stan H.